Adigirl29:

Find the perimeter of the shape. 12.4 units 13.4 units 15.1 units 16.8 units

Adigirl29:

jackiejackiejack:

u just all all the sides up-

Adigirl29:

so its 13.4????

jackiejackiejack:

is that what u got when u added all the sides up

Adigirl29:

I got -3.8

Adigirl29:

so im confused-

jhonyy9:

so to calcule the sides lengths just use the distance formula

Adigirl29:

oh wait so is it 12.4?

jhonyy9:

do you know the distance formula for the distance between two points in a coordinate plane ?

Adigirl29:

not really-

jhonyy9:

d = sqrt [(x_2 -x_1)^2 *(y_2 -y_1)^2]

Adigirl29:

so I have to use this formula when adding up each side?

jhonyy9:

to get the distance between these points in this way to get the length of these sides yes just use thie formula

Adigirl29:

and um how exactly do I use the formula? where do I type in the points of the sides

jhonyy9:

for every side length you need use it separate using the coordinates of these given points

jhonyy9:

for point E x = ? and y = ?

jhonyy9:

for point F the x = ? and y = ? this for example to calcule the length of side EF

Adigirl29:

okay so G is (-1,3) and F is (2,4)

Adigirl29:

so it would be (2-[-1])2+(4-3)2

Adigirl29:

right?

snowflake0531:

H to G: \(2\) G to F: \(\sqrt{10}\) F to E: \(\sqrt{13}\) E to H: \(\sqrt{13}\) just use the pythagorean theorem add them together and find the approximate value